# A Step Toward World Models: A Survey on Robotic Manipulation

### v1

@@ V-H8 Error propagation
Cen et al. [40] indicate that generating multiple actions in sequence leads to performance drop in autoregressive models. The primary reason for this is that pretrained multimodal language models have predominantly been exposed to images and text rather than actions, resulting in limited action generalization capabilities. In autoregressive models where subsequent actions are conditioned on preceding ones, error propagation becomes a critical issue, as the earlier incorrect predictions influence subsequent actions over time. To handle this, Cen et al. [40] propose an attention mask strategy that selectively masks prior actions during the generation of the current action. It enables both future imagination and action generation.
@@ VI Towards Defining Core Components and Capabilities of World Models
From our survey of current approaches, we summarize some potential key components and capabilities that a world model should possess.
@@
Robotic Pushing [30] Laboratory Video, Gripper Pose 59k Robotic Arm Push -
@@
RT-1 [167] Office, Kitchen Video $\sim$ 130k Robot 744 Tasks, e.g., Pick, Move -
@@
Howto100M [168] 12 Env., e.g., Home, Garden Video, Annotation 136M Human 23k Tasks, e.g., Cook, Mark -
@@
Ego4D [169] Dailylife, Outdoor, Indoor Video, Annotation, Audio 3,670 H Human Diverse Tasks, e.g., flip, lift 4D, Egocentric
@@
Ego-Exo4D [28] 123 Natural Scenes Video, Audio, 3D Point Cloud, Annotation 221.26 H Videos, 79M+ 3D Body Poses Human 43 Tasks, 689 Keysteps Multi-view, 4D, Egocentric, Exocentric
@@
ManiFlow-110k [69] - Optical Flow 110k Robot, Human Object Movement 3D
@@
Something-Something V2 [170] Human-object Interaction Video 108+K Human 12 Tasks, e.g., Pour, Drop -
@@
EPIC-KITCHENS [171] Kitchen Video, Action Narrations 432 Sequences/55 H Human 323 Tasks, e.g., Cook, Clean Egocentric
@@
Kinetics-700 [172] Kitchen Video 650K Human 700 Tasks, e.g., Pour, Walk
@@
Bridgedata v2 [173] 24 Env., e.g., Kitchens, Tabletops video 60.1K Traj. Robot 13 Tasks, e.g., Sweep -
@@
GR-2 [58] Table-top Video 40K Traj., Human Diverse tasks/skills, e.g., Pick, Place -
HM3D [174] Photorealistic Indoor RGB-D Video 1K Scenes - - 3D
@@
LanguageTable [175] Simulated & Real-world Tabletop Video 600k Traj. Robot, end-effector Diverse tasks/skills, e.g., Push, Separate -
@@
Bridge Data [176] Toy & Real Kitchens, Toy Sinks Video 7.2K Trajec. Robot 71 Tasks, e.g., Flip, Put -
@@
Matterport3D [177] Building-scale Scenes RGB-D Images 194K+ Tripod-mounted Cameras - Panoramic-view
@@
Multisensory-Universe [18] Indoor Scenes Images, Audio, Tactile, etc 500k Robot 67 Robotic Manip. tasks 3D, Multi-view
@@
Open X-Embodiment [178] 311 Scenes, e.g., Household Video, Depth, 3D Information 1M+ Traj. Robot 527 Skills/160266 Tasks, e.g., Pick, Move 3D
@@
DROID [179] 52 Buildings/564 Scenes RGB Video, Depth, Proprioception, Instructions 76k Traj., 350 H Robot 86 Tasks -
@@
ManiSkill [180] Simulation RGB/RGB-D Video, 3D Point Clouds, Proprioception 4M+ Frames, 2K+ Object Mdels Robot 20 Manip. Tasks 3D
@@
RoboTurk [181] Tabletop RGB & Depth Images, Robot Sensor Data 2K+ Demonstration Frames, 111.25 H Robot Arms Object Search, Tower Creation, Laundry Layout 3D, Long-horizon Object Manip.
@@
AgiBot-World-Beta [112] 106 Scenes Traj. 1M+ Traj., 2.9K H Robot Arms 217 Tasks, 87 Skills Dual-arm, Dexterous hands, Collaboration
@@
3D-EIT [41] Simulation Images, 3D Point Cloud, Annotation 2M Robot, Human Robot Manip., Human-object Interaction 3D
@@
TesserAct [24] Synthetic & Real Scenes RGB, Depth, Normal videos 285K Robot Arms Robotic Manip. 4D, Based on [182, 167, 173, 170]
@@
SynGrasp-1B [136] Tabletop, Simulation Image 1B Frames, 10K+ Objects Robot Grasp -
@@
RH20T-P [183] Tabletop, etc Video 38k Robot 67 Robotic Manip. Tasks Primitive-level, Multi-view
@@
RoboNet [184] Diverse ENV., e.g., arena Video, Traj. 15M Video Frame, 62K Traj. Robot Arm and Gripper Manip. Tasks, e.g., push and pick-and-place -
@@
GenHowTo [165] Kitchens, etc Image Triplets 200k Human Manip., e.g., Cut Tasks -
@@ VII Dataset
There are abundant datasets that facilitate robot learning, including general robotic manipulation datasets [179, 30, 173, 175, 176, 181, 136, 183], Dual-arm robotic manipulation datasets [112], human manipulation datasets [168, 170, 165, 172, 58, 41], combinations of robotic & human manipulation [41], egocentric datasets [169, 28, 171], 3D & 4D datasets [24, 41, 181, 69, 169, 28, 174, 177, 18, 178, 180], multi-view datasets [18, 28, 183] and panoramic-view datasets [177]. A detailed information of them can be found in Fig. II.
Recent efforts in world models for robotic manipulation often have leveraged a large and diverse datasets, e.g., a combinations of different dataset, to be capable of generalizing across tasks and environments. For example, Yang et al. [19] constructed a large-scale natural dataset combining simulated executions and renderings [174, 175], real robot data [176], human activity videos [169, 171, 170], 3D panorama scans [177], and internet text-image data LAION-400M [185]. Bruce et al. [20] combine the RT-1 dataset [167] with real robot grasping data [186]. Wu et al. [6] train the world model based on the combination of the Open X-Embodiment (OXE) dataset [178] and the Something-Something v2 (SSv2) trajectory dataset [170]. Bruce et al. [58] employ a pretraining and fine-tuning strategy. In the pretraining stage, a combination of human demonstration datasets such as Howto100M [168], Ego4D [169], Something-Something V2 [170], EPIC-KITCHENS [171], Kinetics-700 [172], and robot datasets [167, 173]. Fine-tuning data includes 105 table-top tasks via teleoperation covering eight skills (e.g., pick, place). Data augmentation are performed to add new objects or change backgrounds by means of a diffusion model [155] and the Segment Anything Model (SAM) [107], as well as a video generation model [187] to sytheize new videos. Du et al. [59] curated an internet-scale pretraining dataset consisting of 14 million video-text pairs, 60 million image-text pairs [188], LAION-400M [185], and a smaller real-world robotic dataset [176]. Huang et al. [8] constructed multi-anchor view video datasets using public sources including RT-1 [167], Taco-Play [189], ManiSkill [180], BridgeData V2 [173], LanguageTable [175], and RoboTurk [181], augmented with Isaac Sim simulations [190]. [24] construct a 4D embodied video dataset based on previous datasets [182, 167, 173, 170] by measuring depth and normal information.
@@ VIII-A Conclusion
This survey provides a comprehensive evaluation of current approaches to world modeling, examining their relevance for robotic manipulation, underlying architectures, functionalities, key challenges, and proposed solutions. By synthesizing these findings, we offer insights into the nature of real world models and outline the efforts required to advance the field. Our goal is to provide readers with a solid foundation and guide future research directions in world modeling.
@@ VIII-B Future research directions
More Diverse Modalities. The real world contains diverse forms of information, and no single sensory modality can capture its full complexity. This motivates world models capable of perceiving and integrating multiple modalities, including vision, language, action, touch, force, and proprioception, along with their interactions. Early progress has been made in this direction. For example, Hong et al. [18] introduce the Multisensory-Universe dataset, which features interactive scenes enriched with tactile, audio, and temperature signals, generated with the assistance of ChatGPT [191].
Hierarchical World Models. Hierarchical systems play a critical role in building effective world models, as they allow agents to reason across multiple layers of abstraction. However, designing hierarchical models is inherently challenging: separating low-level and high-level dynamics is difficult, and coordinating interactions across layers adds further complexity. While existing studies primarily emphasize model design [56, 11, 45, 80], their validation in complex real-world environments remains limited.
Causality is a fundamental principle for understanding and modeling the world, describing how events or factors influence outcomes and enabling reasoning about future consequences. Causality is the key to world model as it allows agents to interact with the world, which is inline with the human cognition. Richens et al. [192] indicate that learning a causal model is the key to ensure the generalization ability to new domains. Wang et al. [193, 194] learn a causal dynamics model by removing unnecessary dependencies for tasks, which however are constrained to specific tasks. Gupta et al. [195] argue that conventional theory-driven approaches to causal modeling, such as those in [196, 197], are insufficient for world models that aim for generalizable understanding. These methods rely on predefined variables and case-specific theoretical properties. In the real world, sensory inputs are complex, often unstructured, and key theoretical properties,such as identifiability, may not hold.
Resource-Constrained Deployment. Current world models, particularly those based on video generation, are computationally intensive and contain hundreds of millions of parameters, which limits their feasibility for real-world robotic deployment and on-device inference. To enable practical applications, designing lightweight and efficient world models has become increasingly important. Quantization and model compression techniques offer promising directions for reducing memory and computational costs, and have been extensively explored in related domains [198, 199, 200, 201], providing both direct solutions and inspiration for future lightweight world model architectures.
@@
Furthermore, deep models are known to be vulnerable to adversarial attacks, which can compromise performance by introducing imperceptible perturbations to inputs [202, 203], modifying model parameters [204, 205], or even exploiting hardware-level weaknesses [206, 207]. These vulnerabilities raise serious concerns regarding the security and reliability of world models, especially when deployed in safety-critical domains. To date, systematic studies on the robustness and security of world models remain limited, underscoring an urgent need for dedicated research into adversarial resilience, trustworthy deployment, and secure model adaptation.
Evaluation Protocols. Current evaluation practices for world models are fragmented and only loosely aligned with their intended capabilities, often relying on task-specific or proxy metrics and partial human validation [57]. There is a pressing need for standardized benchmarks and unified evaluation frameworks that can comprehensively assess world model competence across multiple dimensions, including visual fidelity, policy success, causal consistency, physical plausibility, generalization, and long-horizon reasoning.
Beyond Human Intelligence. Insights from human cognition have profoundly influenced the design of robotic and world modeling systems. However, the completeness of the world extends beyond human cognition, which is bounded by partial observation, finite memory, limited attention, and inherent heuristic biases. World models are therefore expected to transcend human cognitive bounds, providing a deeper and more systematic understanding of complex environments.

### v2

@@ V-H8 Error Propagation
In autoregressive models, subsequent actions are generated based on previous predictions, leading to error propagation over time. To handle this, Cen et al. [cen2025worldvla] propose an attention mask strategy that selectively masks prior actions during the generation of the current action. It enables both future imagination and action generation.
@@ VI Towards Defining Core Components and Capabilities of World Models
From our survey of current approaches, we summarize some potential key components and capabilities that a world model should possess. Future research may identify additional dimensions necessary for comprehensive world modeling.
Figure 4: Potential Core Components and Capabilities of World Models.
@@
Robotic Pushing [ebert2018robustness] Laboratory Video, Gripper Pose 59k Robotic Arm Push -
@@
RT-1 [brohan2023rt] Office, Kitchen Video $\sim$ 130k Robot 744 Tasks, e.g., Pick, Move -
@@
Howto100M [miech2019howto100m] 12 Env., e.g., Home, Garden Video, Annotation 136M Human 23k Tasks, e.g., Cook, Mark -
@@
Ego4D [grauman2022ego4d] Dailylife, Outdoor, Indoor Video, Annotation, Audio 3,670 H Human Diverse Tasks, e.g., flip, lift 4D, Egocentric
@@
Ego-Exo4D [grauman2024ego] 123 Natural Scenes Video, Audio, 3D Point Cloud, Annotation 221.26 H Videos, 79M+ 3D Body Poses Human 43 Tasks, 689 Keysteps Multi-view, 4D, Egocentric, Exocentric
@@
ManiFlow-110k [zhi20253dflowaction] - Optical Flow 110k Robot, Human Object Movement 3D
@@
Something-Something V2 [goyal2017something] Human-object Interaction Video 108+K Human 12 Tasks, e.g., Pour, Drop -
@@
EPIC-KITCHENS [damen2018scaling] Kitchen Video, Action Narrations 432 Sequences/55 H Human 323 Tasks, e.g., Cook, Clean Egocentric
@@
Kinetics-700 [carreira2019short] Kitchen Video 650K Human 700 Tasks, e.g., Pour, Walk
@@
Bridgedata v2 [walke2023bridgedata] 24 Env., e.g., Kitchens, Tabletops video 60.1K Traj. Robot 13 Tasks, e.g., Sweep -
@@
GR-2 [cheang2024gr] Table-top Video 40K Traj., Human Diverse tasks/skills, e.g., Pick, Place -
HM3D [ramakrishnan22021habitat] Photorealistic Indoor RGB-D Video 1K Scenes - - 3D
@@
LanguageTable [lynch2023interactive] Simulated & Real-world Tabletop Video 600k Traj. Robot, end-effector Diverse tasks/skills, e.g., Push, Separate -
@@
Bridge Data [ebert2021bridge] Toy & Real Kitchens, Toy Sinks Video 7.2K Trajec. Robot 71 Tasks, e.g., Flip, Put -
@@
Matterport3D [chang2017matterport3d] Building-scale Scenes RGB-D Images 194K+ Tripod-mounted Cameras - Panoramic-view
@@
Multisensory-Universe [hong2024multiply] Indoor Scenes Images, Audio, Tactile, etc 500k Robot 67 Robotic Manip. tasks 3D, Multi-view
@@
Open X-Embodiment [o2024open] 311 Scenes, e.g., Household Video, Depth, 3D Information 1M+ Traj. Robot 527 Skills/160266 Tasks, e.g., Pick, Move 3D
@@
DROID [khazatsky2024droid] 52 Buildings/564 Scenes RGB Video, Depth, Proprioception, Instructions 76k Traj., 350 H Robot 86 Tasks -
@@
ManiSkill [gu2023maniskill2] Simulation RGB/RGB-D Video, 3D Point Clouds, Proprioception 4M+ Frames, 2K+ Object Mdels Robot 20 Manip. Tasks 3D
@@
RoboTurk [mandlekar2019scaling] Tabletop RGB & Depth Images, Robot Sensor Data 2K+ Demonstration Frames, 111.25 H Robot Arms Object Search, Tower Creation, Laundry Layout 3D, Long-horizon Object Manip.
@@
AgiBot-World-Beta [bu2025agibot] 106 Scenes Traj. 1M+ Traj., 2.9K H Robot Arms 217 Tasks, 87 Skills Dual-arm, Dexterous hands, Collaboration
@@
3D-EIT [zhen20243d] Simulation Images, 3D Point Cloud, Annotation 2M Robot, Human Robot Manip., Human-object Interaction 3D
@@
TesserAct [zhen2025tesseract] Synthetic & Real Scenes RGB, Depth, Normal videos 285K Robot Arms Robotic Manip. 4D, Based on [james2020rlbench, brohan2023rt, walke2023bridgedata, goyal2017something]
@@
SynGrasp-1B [deng2025graspvla] Tabletop, Simulation Image 1B Frames, 10K+ Objects Robot Grasp -
@@
RH20T-P [chen2024rh20t] Tabletop, etc Video 38k Robot 67 Robotic Manip. Tasks Primitive-level, Multi-view
@@
RoboNet [dasari2020robonet] Diverse ENV., e.g., arena Video, Traj. 15M Video Frame, 62K Traj. Robot Arm and Gripper Manip. Tasks, e.g., push and pick-and-place -
@@
GenHowTo [souvcek2024genhowto] Kitchens, etc Image Triplets 200k Human Manip., e.g., Cut Tasks -
@@ VII Dataset
There are abundant datasets that facilitate robot learning, including general robotic manipulation datasets [khazatsky2024droid, ebert2018robustness, walke2023bridgedata, lynch2023interactive, ebert2021bridge, mandlekar2019scaling, deng2025graspvla, chen2024rh20t], dual-arm robotic manipulation datasets [bu2025agibot], human manipulation datasets [miech2019howto100m, goyal2017something, souvcek2024genhowto, carreira2019short, cheang2024gr, zhen20243d], combinations of robotic & human manipulation [zhen20243d], egocentric datasets [grauman2022ego4d, grauman2024ego, damen2018scaling], 3D & 4D datasets [zhen2025tesseract, zhen20243d, mandlekar2019scaling, zhi20253dflowaction, grauman2022ego4d, grauman2024ego, ramakrishnan22021habitat, chang2017matterport3d, hong2024multiply, o2024open, gu2023maniskill2], multi-view datasets [hong2024multiply, grauman2024ego, chen2024rh20t] and panoramic-view datasets [chang2017matterport3d]. A detailed information can be found in Table. II.
Recent efforts in world models for robotic manipulation have leveraged large and diverse datasets, often combining multiple sources, to enable generalization across tasks and environments. For example, Yang et al. [yang2023learning] construct a large-scale natural dataset combining simulated executions and renderings [ramakrishnan22021habitat, lynch2023interactive], real robot data [ebert2021bridge], human activity videos [grauman2022ego4d, damen2018scaling, goyal2017something], 3D panorama scans [chang2017matterport3d], and internet text-image data (i.e., LAION-400M [schuhmann2021laion]). Bruce et al. [bruce2024genie] combine the RT-1 dataset [brohan2023rt] with real robot grasping data [kalashnikov2018scalable]. Wu et al. [wu2024ivideogpt] train the world model based on the combination of the Open X-Embodiment (OXE) dataset [o2024open] and the Something-Something v2 trajectory dataset [goyal2017something]. Bruce et al. [cheang2024gr] employ a pretraining and fine-tuning strategy. In the pretraining stage, a combination of human demonstration datasets such as Howto100M [miech2019howto100m], Ego4D [grauman2022ego4d], Something-Something V2 [goyal2017something], EPIC-KITCHENS [damen2018scaling], Kinetics-700 [carreira2019short], and robot datasets [brohan2023rt, walke2023bridgedata]. Fine-tuning data includes 105 table-top tasks via teleoperation covering eight skills (e.g., pick, place). Data augmentations are performed to add new objects or change backgrounds by means of a diffusion model [ho2020denoising] and the Segment Anything Model (SAM) [kirillov2023segment], as well as a video generation model [ma2025latte] to synthesize new videos. Du et al. [du2023learning] curate an internet-scale pretraining dataset consisting of 14 million video-text pairs, 60 million image-text pairs [ho2022imagen], LAION-400M [schuhmann2021laion], and a smaller real-world robotic dataset [ebert2021bridge]. Huang et al. [huang2025enerverse] construct multi-anchor view video datasets using public sources including RT-1 [brohan2023rt], Taco-Play [rosete2023latent], ManiSkill [gu2023maniskill2], BridgeData V2 [walke2023bridgedata], LanguageTable [lynch2023interactive], and RoboTurk [mandlekar2019scaling], augmented with Isaac Sim simulations [mittal2023orbit]. [zhen2025tesseract] construct a 4D embodied video dataset based on previous datasets [james2020rlbench, brohan2023rt, walke2023bridgedata, goyal2017something] by measuring depth and normal information.
@@ VIII-A Conclusion
This survey provides a comprehensive evaluation of current approaches to world modeling, examining their relevance for robotic manipulation, underlying architectures, functionalities, key challenges, and proposed solutions. By synthesizing these findings, we offer insights into the nature of fully realized world models and outline the efforts required to advance the field. Our goal is to provide readers with a solid foundation and guide future research directions in world modeling.
@@ VIII-B Future research directions
More Diverse Modalities. The real world contains diverse forms of information, and no single sensory modality can capture its full complexity. This motivates world models capable of perceiving and integrating multiple modalities, including vision, language, action, touch, force, and proprioception, along with their interactions. Early progress has been made in this direction. For example, Hong et al. [hong2024multiply] introduce the Multisensory-Universe dataset, which features interactive scenes enriched with tactile, audio, and temperature signals, generated with the assistance of ChatGPT [achiam2023gpt].
Hierarchical World Models. Hierarchical systems play a critical role in building effective world models, as they allow agents to reason across multiple layers of abstraction. However, designing hierarchical models is inherently challenging: separating low-level and high-level dynamics is difficult, and coordinating interactions across layers adds further complexity. While existing studies primarily emphasize model design [gumbsch2023learning, lecun2022path, wang2025dmwm, xing2025critiques], their validation in complex real-world environments remains limited.
Causality is a fundamental principle for understanding and modeling the world, describing how events or factors influence outcomes and enabling reasoning about future consequences. It is the key to the world model and agents to allow them to interact with the world, which is inline with the human cognition. Richens et al. [richens2024robust] indicate that learning a causal model is the key to ensure the generalization ability to new domains. Wang et al. [wang2022causal, tomar2021model] learn a causal dynamics model by removing unnecessary dependencies for tasks, which however are constrained to specific tasks. Gupta et al. [gupta2024essential] argue that conventional theory-driven approaches to causal modeling, such as those in [stuart2010matching, chernozhukov2018double], are insufficient for world models that aim for generalizable understanding. These methods rely on predefined variables and case-specific theoretical properties. In the real world, sensory inputs are complex, often unstructured, and key theoretical properties, such as identifiability, may not hold.
Resource-Constrained Deployment. Current world models, particularly those based on video generation, are computationally intensive and contain hundreds of millions of parameters, which limits their feasibility for real-world robotic deployment and on-device inference. To enable practical applications, designing lightweight and efficient world models has become increasingly important. Quantization and model compression techniques offer promising directions for reducing memory and computational costs, and have been extensively explored in related domains [polino2018model, gholami2022survey, shang2023post, li2021lightweight], providing both direct solutions and inspiration for future lightweight world model architectures.
@@
Furthermore, deep models are known to be vulnerable to adversarial attacks, which can compromise performance by introducing imperceptible perturbations to inputs [szegedy2013intriguing, zhang2024universal], modifying model parameters [ren2023dimension, park2022blurs], or even exploiting hardware-level weaknesses [cojocar2020we, jattke2024zenhammer]. These vulnerabilities raise serious concerns regarding the security and reliability of world models, especially when deployed in safety-critical domains. To date, systematic studies on the robustness and security of world models remain limited, underscoring an urgent need for dedicated research into adversarial resilience, trustworthy deployment, and secure model adaptation.
Evaluation Protocols. Current evaluation practices for world models are fragmented and only loosely aligned with their intended capabilities, often relying on task-specific or proxy metrics and partial human validation [liao2025genie]. There is a pressing need for standardized benchmarks and unified evaluation frameworks that can comprehensively assess world model competence across multiple dimensions, including visual fidelity, policy success, causal consistency, physical plausibility, generalization, and long-horizon reasoning.
Beyond Human Intelligence. Insights from human cognition have profoundly influenced the design of robotic and world modeling systems [lecun2022path, gumbsch2023learning, bjorck2025gr00t, wang2025dmwm]. However, the completeness of the world extends beyond human cognition, which is bounded by partial observation, finite memory, limited attention, and inherent heuristic biases. World models are therefore expected to transcend human cognitive bounds, providing a deeper and more systematic understanding of complex environments.